# Lampr: Boosting the Effectiveness of Language-Generic Program Reduction via Large Language Models

### 20 Dec 2023 v1

@@ 4.3. RQ2: Efficiency
In this research question, we measure the time consumed by each technique on three benchmarks. Shorter time indicates higher efficiency. Table 2 shows the results. Since both Vulcan and Lampr perform reduction on top of Perses’s results, it is impossible for these two algorithms to take less time than Perses. Besides, as a highly efficient tree-based reduction algorithm, Perses is generally faster than C-Reduce. Therefore, we focus on the comparison among Vulcan, C-Reduce and Lampr.
| Benchmark | Perses | Vulcan | C-Reduce | Lampr | LPR+Vulcan | |||
@@
In the Benchmark-C, compared to Vulcan and Lampr, C-Reduce generally has a shorter reduction time. This is expected, as C-Reduce’s transformations are specifically designed for C languages, whereas Vulcan approaches the problem in a more unguided and brute-force manner, and Lampr’s attempts are not specifically designed for C and rely more on the efficiency of LLMs. On average, Lampr takes 1:54:54, which is % shorter than 2:08:46 of Vulcan. However, in terms of the average percentage difference in time, Lampr requires 45.83% more time compared to Vulcan in Benchmark-C. The main reason of such a result is that Lampr is more efficient than Vulcan when the program is large and complex, as the transformations it performs are aware of the semantics and have a higher success rate in reducing the program. However, when the program is small and simple, Vulcan can finish quickly since its search space becomes considerably small, but for Lampr, the time consumed by the LLM is not decreased significantly and becomes dominant, which makes Lampr less efficient than Vulcan in these benchmarks. On Benchmark-Rust, The results indicate a similar trend, i.e., Lampr tends to be more efficient when reducing complex programs while Vulcan is more efficient in small and simple benchmarks. If we only keep benchmarks where both tools take longer than one hour, Lampr requires 4.15% and 21.69% less time compared to Vulcan.
In our analysis of Benchmark-Rust, we observed that both Lampr and Vulcan can consume an extreme long time on certain benchmarks. For instance, Vulcan requires 20 hours for Rust-99830, while Lampr takes a similar duration on Rust-66851 in a specific run. This extensive time consumption is often due to the frequent invocation of Perses, which only achieves marginal progress with each transformation. The prolonged duration of Perses is primarily attributed to the strict syntax of the Rust language. Considering that program reduction is an NP-complete problem, this inefficiency might be optimized in future, whereas it cannot be completely eliminated.
After further in-depth analysis, another interesting fact emerges. On the three benchmark suites, the average time taken by Lampr is 1.915, 1.839, and 0.174 hours, respectively. However, within these durations, the time spent waiting for the LLM responses accounts for %, %, and %. This experiment involves invoking the OpenAI API, and might have been limited by high user traffic and few computational resources allocated. Considering the ongoing advancements in LLMs technology, we believe that Lampr’s efficiency will be substantially improved in the future.
@@ 5.2. The Impact of Temperature
| Perses | Vulcan | Lampr with gpt-3.5-turbo-0613 | |||||
| Mean | 161.4 | 102.8 | |||||
As shown in Table 3, the performances under most of the temperatures are similar, while the exception is t=0, with the average size worse than others. According to the documentation of temperature (ope, 2023c), t=0 will actually use a small threshold above 0. Our speculation is that a low temperature restrains the diversity of outputs, impeding Lampr exploring local minima in different runs, which is helpful in program reduction tasks.
Given that program reduction is an NP-complete problem, randomness of LLMs has its advantages and drawbacks. Take the results of RQ1 in Table 1 as an example. On the one hand, randomness allows Lampr to explore more distinct local minima, and sometimes generates smaller programs than C-Reduce in five repeated experiments on each benchmark. On the other hand, randomness of LLMs introduces variability. While the standard deviation remains below 10 in most benchmarks, it may significantly increase in certain benchmarks, such as in GCC-59903. This variation can be attributed to the high complexity of the given program. In such scenarios, LLMs might not consistently execute accurate transformations, resulting in a range of local minima and divergent results.
@@ 5.3.2. Threats to External Validity
One threat to external validity is the generality of Lampr across languages. We evaluate Lampr on three prevalent programming languages, namely, C, Rust and JavaScript. We believe these benchmark suites have demonstrated Lampr’s high generality on diverse programs. However, without further experiments on a particular language, we still cannot ensure the effectiveness and efficiency of Lampr on that language.
@@ 6.1. Program Reduction
Distinct from prior work, Lampr synergistically combines LLMs and language-generic reduction tool to harness the advantages of both. Language-generic reducers stand out as their remarkable generality across multiple languages, while LLMs excel in further refining the programs with the domain knowledge on certain languages learned from large training sets. Language-specific tools typically demand considerable human effort to design and implement feature-related transformations for reduction, while Lampr requires only a few lines of natural languages prompts, significantly reducing the effort involved.
@@ 6.2. LLMs for Software Engineering
Similar to these studies, our approach Lampr leverages LLMs for a software engineering task, i.e., program reduction. Lampr harnesses the comprehension and generation capabilities of LLMs to refine the results of program reduction. However, our work distinguishes itself in the nature of the programs processed by LLMs. In related research, programs are typically logical and goal-oriented, often designed to fulfill a specific purpose. In contrast, the programs involved in our program reduction task are random, chaotic, and lack a clear objective. Consequently, our research sheds light on the performance of LLMs when dealing with programs that do not have an easily discernible purpose.
@@ 7. Conclusion
This paper proposes LLMs-aided program reduction (Lampr), which is the first approach that leverages LLMs for the program reduction task to the best of our knowledge. By combining the strength of LLMs and existing language-generic program reduction techniques, Lampr can perform language-specific transformations to effectively reduce the program while being language-generic (i.e., can be easily applied to a wide range of languages). The evaluation shows that in 50 benchmarks across three programming languages. Lampr significantly outperforms Vulcan. Specifically, Lampr produces 24.93%, 4.47%, and 11.71% smaller programs on C, Rust, and JavaScript, respectively. Meanwhile, The evaluation also demonstrates that Lampr complements Vulcan to some extent. By reducing the outputs of Lampr with Vulcan, we attained results that have similar sizes to those of C-Reduce in Benchmark-C. In terms of efficiency, Lampr excels in reducing complex programs, and takes %, %, % less time than Vulcan to finish all the benchmarks, respectively.

### 02 Mar 2024 v2

@@ 4.3. RQ2: Efficiency
In this research question, we measure the time consumed by each technique on three benchmarks. Shorter time indicates higher efficiency. Table 2 shows the results. Since both Vulcan and LPR perform reduction on top of Perses’s results, it is impossible for these two algorithms to take less time than Perses. Besides, as a highly efficient tree-based reduction algorithm, Perses is generally faster than C-Reduce. Therefore, we focus on the comparison among Vulcan, C-Reduce and LPR.
Benchmark | Perses | Vulcan | C-Reduce | LPR | LPR+Vulcan | |||
@@
In the Benchmark-C, compared to Vulcan and LPR, C-Reduce generally has a shorter reduction time. This is expected, as C-Reduce’s transformations are specifically designed for C languages, whereas Vulcan approaches the problem in a more unguided and brute-force manner, and LPR’s attempts are not specifically designed for C and rely more on the efficiency of LLMs. On average, LPR takes 1:54:54, which is % shorter than 2:08:46 of Vulcan. However, in terms of the average percentage difference in time, LPR requires 45.83% more time compared to Vulcan in Benchmark-C. The main reason of such a result is that LPR is more efficient than Vulcan when the program is large and complex, as the transformations it performs are aware of the semantics and have a higher success rate in reducing the program. However, when the program is small and simple, Vulcan can finish quickly since its search space becomes considerably small, but for LPR, the time consumed by the LLM is not decreased significantly and becomes dominant, which makes LPR less efficient than Vulcan in these benchmarks. On Benchmark-Rust, The results indicate a similar trend, i.e., LPR tends to be more efficient when reducing complex programs while Vulcan is more efficient in small and simple benchmarks. If we only keep benchmarks where both tools take longer than one hour, LPR requires 4.15% and 21.69% less time compared to Vulcan.
In our analysis of Benchmark-Rust, we observed that both LPR and Vulcan can consume an extreme long time on certain benchmarks. For instance, Vulcan requires 20 hours for Rust-99830, while LPR takes a similar duration on Rust-66851 in a specific run. This extensive time consumption is often due to the frequent invocation of Perses, which only achieves marginal progress with each transformation. The prolonged duration of Perses is primarily attributed to the strict syntax of the Rust language. Considering that program reduction is an NP-complete problem, this inefficiency might be optimized in future, whereas it cannot be completely eliminated.
After further in-depth analysis, another interesting fact emerges. On the three benchmark suites, the average time taken by LPR is 1.915, 1.839, and 0.174 hours, respectively. However, within these durations, the time spent waiting for the LLM responses accounts for %, %, and %. This experiment involves invoking the OpenAI API, and might have been limited by high user traffic and few computational resources allocated. Considering the ongoing advancements in LLMs technology, we believe that LPR’s efficiency will be substantially improved in the future.
@@ 5.2. The Impact of Temperature
Perses | Vulcan | LPR with gpt-3.5-turbo-0613 | |||||
Mean | 161.4 | 102.8 |
As shown in Table 3, the performances under most of the temperatures are similar, while the exception is t=0, with the average size worse than others. According to the documentation of temperature (ope, 2023c), t=0 will actually use a small threshold above 0. Our speculation is that a low temperature restrains the diversity of outputs, impeding LPR exploring local minima in different runs, which is helpful in program reduction tasks.
Given that program reduction is an NP-complete problem, randomness of LLMs has its advantages and drawbacks. Take the results of RQ1 in Table 1 as an example. On the one hand, randomness allows LPR to explore more distinct local minima, and sometimes generates smaller programs than C-Reduce in five repeated experiments on each benchmark. On the other hand, randomness of LLMs introduces variability. While the standard deviation remains below 10 in most benchmarks, it may significantly increase in certain benchmarks, such as in GCC-59903. This variation can be attributed to the high complexity of the given program. In such scenarios, LLMs might not consistently execute accurate transformations, resulting in a range of local minima and divergent results.
@@ 5.3.2. Threats to External Validity
One threat to external validity is the generality of LPR across languages. We evaluate LPR on three prevalent programming languages, namely, C, Rust and JavaScript. We believe these benchmark suites have demonstrated LPR’s high generality on diverse programs. However, without further experiments on a particular language, we still cannot ensure the effectiveness and efficiency of LPR on that language.
@@ 6.1. Program Reduction
Distinct from prior work, LPR synergistically combines LLMs and language-generic reduction tool to harness the advantages of both. Language-generic reducers stand out as their remarkable generality across multiple languages, while LLMs excel in further refining the programs with the domain knowledge on certain languages learned from large training sets. Language-specific tools typically demand considerable human effort to design and implement feature-related transformations for reduction, while LPR requires only a few lines of natural languages prompts, significantly reducing the effort involved.
@@ 6.2. LLMs for Software Engineering
Similar to these studies, our approach LPR leverages LLMs for a software engineering task, i.e., program reduction. LPR harnesses the comprehension and generation capabilities of LLMs to refine the results of program reduction. However, our work distinguishes itself in the nature of the programs processed by LLMs. In related research, programs are typically logical and goal-oriented, often designed to fulfill a specific purpose. In contrast, the programs involved in our program reduction task are random, chaotic, and lack a clear objective. Consequently, our research sheds light on the performance of LLMs when dealing with programs that do not have an easily discernible purpose.
@@ 7. Conclusion
This paper proposes LLMs-aided program reduction (LPR), which is the first approach that leverages LLMs for the program reduction task to the best of our knowledge. By combining the strength of LLMs and existing language-generic program reduction techniques, LPR can perform language-specific transformations to effectively reduce the program while being language-generic (i.e., can be easily applied to a wide range of languages). The evaluation shows that in 50 benchmarks across three programming languages. LPR significantly outperforms Vulcan. Specifically, LPR produces 24.93%, 4.47%, and 11.71% smaller programs on C, Rust, and JavaScript, respectively. Meanwhile, The evaluation also demonstrates that LPR complements Vulcan to some extent. By reducing the outputs of LPR with Vulcan, we attained results that have similar sizes to those of C-Reduce in Benchmark-C. In terms of efficiency, LPR excels in reducing complex programs, and takes %, %, % less time than Vulcan to finish all the benchmarks, respectively.